OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-5, 1)with a slope of . Answer y - 1 = (x + 5) y + 1 = (x - 5) y - 5 = (x + 1) y + 5 = -(x - 1)

jimthompson5910 (jim_thompson5910):

Slope of 1 or -1?

OpenStudy (anonymous):

slope of 3/2

jimthompson5910 (jim_thompson5910):

\[\Large y - y_{1} = m(x-x_{1})\] \[\Large y - 1 = \frac{3}{2}(x-(-5))\] \[\Large y - 1 = \frac{3}{2}(x+5)\]